OpenStudy (anonymous):

How do i define the variables of y=.3x+12

OpenStudy (mertsj):

x is the independent variable and y is the dependent variable.

OpenStudy (anonymous):

but how do i solve them

OpenStudy (anonymous):

???

OpenStudy (anonymous):

heeeelp!

OpenStudy (mertsj):

You cannot solve 1 equation with 2 variables unless you have more information. You obviously haven't posted the entire problem.

OpenStudy (anonymous):

It says i have to define the variables of y= .3x+12 and y=.15+20

OpenStudy (mertsj):

From your second equation, y = .15 +20 Therefore y = 20.15 In the first equation replace y with 20.15 and solve for x

OpenStudy (anonymous):

how do i identify the slope and x and y intercept

OpenStudy (mertsj):

y=mx+b The slope is m and the y intercept is b

OpenStudy (anonymous):

were did you get m and b from